OpenStudy (anonymous):

Help plz! 2^x • 5^X=10^20 WHAT IS X ?

OpenStudy (anonymous):

help meh

OpenStudy (fibonaccichick666):

how can you bring the exponent down?

OpenStudy (anonymous):

divide ?

OpenStudy (anonymous):

i wouldn't bring it down \[2^x\times 5^x=(2\times 5)^x=10^x\]

OpenStudy (fibonaccichick666):

no, you can't divide, but satellite's was my second step

OpenStudy (anonymous):

ok ?

OpenStudy (anonymous):

so you have \[10^x=10^{20}\] and so \(x\) should be easy enough to find

OpenStudy (fibonaccichick666):

May approach wasn't nearly that simple. Nice catch @satellite73

OpenStudy (anonymous):

wait why from 10^20 turn into 20^x ?

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

nevermind got it !

OpenStudy (anonymous):

there is no \(20^x\) in the question the left hand side of the equal sign is \(10^x\) and the right hand side is \(10^{20}\)

OpenStudy (fibonaccichick666):

That was sat's steps

OpenStudy (fibonaccichick666):

Do you see what he did?

OpenStudy (anonymous):

yep

OpenStudy (fibonaccichick666):

whoops wait a sec

OpenStudy (fibonaccichick666):

\[2^x×5^x=10^{20}\] \[(2×5)^x=10^{20}\] \[10^x=10^{20}\]

OpenStudy (anonymous):

don't you cross 10 out becuase they are the same ?

OpenStudy (anonymous):

so it would be x = 20

OpenStudy (fibonaccichick666):

no, not exactly

OpenStudy (fibonaccichick666):

yes x=20 but you can't just cross out the 10

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (fibonaccichick666):

You use logarithms